OpenStudy (anonymous):

cos(pi/11)+cos(3pi/11)+cos(5pi/11) + cos(7pi/11) + cos(9pi/11)

OpenStudy (anonymous):

It is 1/2

OpenStudy (anonymous):

but my answer is not matchig the text

OpenStudy (anonymous):

USE COSINE SERIES FORMULAE

OpenStudy (anonymous):

@rational

OpenStudy (ikram002p):

It should bezero , right ?

OpenStudy (anonymous):

it is 0.5

OpenStudy (ikram002p):

It should bezero , right ?

OpenStudy (anonymous):

does anyone know the cosine series formula

OpenStudy (anonymous):

It is to be solved by that method only

OpenStudy (ikram002p):

Will try ltr my phone acting weird lol i had to switch to laptob to use os

OpenStudy (rational):

\[\large \begin{align} e^{i \pi/11} + e^{i3\pi/11} + \cdots + e^{i9\pi/11} &= \dfrac{e^{i \pi/11} \left(e^{i 10\pi/11} - 1\right)}{e^{i 2\pi/11} -1} \\~\\ &= \dfrac{e^{i\pi} - e^{i \pi/11} }{(e^{i \pi/11} -1)(e^{i \pi/11} +1)}\\~\\&= \dfrac{-1-e^{i \pi/11}}{(e^{i \pi/11} -1)(e^{i \pi/11} +1)}\\~\\&= \dfrac{1}{1-e^{i\pi/11}}\end{align}\]

OpenStudy (anonymous):

not that

OpenStudy (rational):

is that the series formula you talking about ?

OpenStudy (anonymous):

:)

OpenStudy (rational):

a = pi/11 d = 2pi/11

OpenStudy (rational):

just plugin and evaluate ?

OpenStudy (anonymous):

i m not getting a answer actually

OpenStudy (rational):

\[\large \dfrac{\sin \left(5\times \pi/11\right)}{\sin \left(\pi/11\right)}\times \cos\left(\dfrac{2\pi/11 + (5-1)2\pi/11}{2}\right)\]

OpenStudy (rational):

\[\large \dfrac{\sin \left(5\pi/11\right)}{\sin \left(\pi/11\right)}\times \cos\left(5\pi/11\right)\]

OpenStudy (rational):

use sinAcosB formula

OpenStudy (anonymous):

that would be sin 10 pi /11 right

OpenStudy (rational):

use sin2A formula

OpenStudy (rational):

yes and with a 1/2

OpenStudy (rational):

\[\large \dfrac{\sin \left(10\pi/11\right)}{2\sin \left(\pi/11\right)}\]

OpenStudy (anonymous):

no wait 1 min

OpenStudy (anonymous):

difference is 2pi/11 right

OpenStudy (rational):

yes d = 2pi/11

OpenStudy (rational):

n = number of terms in the series = 5

OpenStudy (anonymous):

ok got it sry , thank you

OpenStudy (anonymous):

I am sorry

OpenStudy (anonymous):